Determine the point of intersecting of the 2 lines



⟨x,y,z⟩=⟨4,2,4⟩+⟨2,2,1⟩t and


⟨x,y,z⟩=⟨-5,2,-8 ⟩+⟨ 20,8,20⟩s


( Blank 1. Calculate the answer by read surrounding text.


, Blank 2. Calculate the answer by read surrounding text.


, Blank 3. Calculate the answer by read surrounding text.


)


1
Expert's answer
2022-03-14T14:58:36-0400

"4+2t=-5+20s" ......(i)

"2+2t=2+8s.....(ii)"


"(i)-(ii)"


"2=-7+12s"

"s=\frac{3}{4}"


From "(i)"

"2t=-9+20(\frac{3}{4})"

"t=3"


Point of intersection


"4+2(3)=10\\
2+2(3)=8\\4+3=7"


"(10,8,7)"